# The Epistemic Support-Point Filter: Jaynesian Maximum Entropy Meets Popperian Falsification A Possibilistic Minimax-Entropy Optimality Proof

### 10 Mar 2026 v1

@@ 3.3 Lemma 3: Minimax Entropy and Evidence-Only Selection
Let $\{\chi^{(i)}\}_{i=1}^{M}$ be the predicted support with whitened squared innovations $q^{(i)}_{k}$ as in Definition 2.8, and let $N_{\mathrm{target}}\leq M$ be a fixed survivor count. Consider all pairs $(J,\rho)$ where $J\subseteq\{1,\ldots,M\}$ with $|J|=N_{\mathrm{target}}$ and $\rho:J\to(0,1]$ with $\max_{i\in J}\rho^{(i)}=1$, within the evidence-only comparison class: all $(J,\rho)$ whose selection and ranking depend only on $q^{(i)}_{k}$, not on $\pi^{(i)}_{k|k-1}$.
@@
$H(J,\rho)=\int_{0}^{1}\!\log V^{J}_{\alpha}\,d\alpha,\qquad V^{J}_{\alpha}=c_{n}\bigl(\det\Pi^{J}_{\alpha}\bigr)^{1/2},$ (12)
@@
$J^{\star}=\bigl\{i:q^{(i)}_{k}\text{ is among the }N_{\mathrm{target}}\text{ smallest}\bigr\},\qquad\rho^{\star}(i)=\mathrm{Comp}^{(i)}_{k}=e^{-q^{(i)}_{k}/2},$ (13)
@@
Equivalently, $J^{\star}$ is the unique minimizer of $\log\det\bigl(\mathrm{MVEE}(\{\chi^{(i)}:i\in J\})\bigr)$ within the evidence-only class.
@@
We show $(J^{\star},\rho^{\star})$ minimizes $H(J,\rho)$ by showing it simultaneously minimizes $V^{J}_{\alpha}$ for every $\alpha\in(0,1]$.
Step 1: $\alpha$-cut structure under $\rho^{\star}$. Under $\rho^{\star}(i)=e^{-q^{(i)}_{k}/2}$, higher possibility corresponds to lower $q^{(i)}_{k}$. For any threshold $\alpha$,
$C_{\alpha}(J^{\star},\rho^{\star})=\bigl\{i\in J^{\star}:e^{-q^{(i)}_{k}/2}\geq\alpha\bigr\}=\bigl\{i\in J^{\star}:q^{(i)}_{k}\leq-2\log\alpha\bigr\}.$ (14)
At $\alpha\to 0$ this is the full survivor set $J^{\star}$.
@@
$q^{(i)}_{k}=\left\|L_{e}^{-1}H(\chi^{(i)}-\chi^{\star})\right\|_{2}^{2}+\mathrm{const},$ (15)
@@
where $\chi^{\star}=H^{\dagger}(y_{k}-b)$ is the evidence-consistent locus. Hence $q^{(i)}_{k}$ measures each hypothesis’s distance to $\chi^{\star}$ in the innovation-whitened metric.
@@
Among all $N_{\mathrm{target}}$-element subsets of $\{\chi^{(i)}\}_{i=1}^{M}$, the set geometrically closest to $\chi^{\star}$ has the smallest MVEE. This follows from the convexity of $\log\det(\cdot)$ on the positive-definite cone [10]: for any set $J$ containing a hypothesis $j$ with $q^{(j)}_{k}>q^{(i^{*})}_{k}$ for some $i^{*}\notin J$, replacing $j$ with $i^{*}$ (moving the set closer to $\chi^{\star}$) does not increase the MVEE determinant, and strictly decreases it for generic (non-degenerate) point clouds.
@@
Therefore $J^{\star}$ minimizes $\log\det(\mathrm{MVEE}(\{\chi^{(i)}:i\in J\}))$ among all $N_{\mathrm{target}}$-element subsets. The same reasoning applies at every $\alpha$-level: for each cardinality $k_{\alpha}$, the $k_{\alpha}$ smallest-$q$ hypotheses minimize $V^{J}_{\alpha}$. Hence $C_{\alpha}(J^{\star},\rho^{\star})$ minimizes $V^{J}_{\alpha}$ for every $\alpha\in(0,1]$.
@@
Step 3: Pointwise minimum volume implies minimum $H_{\pi}$. Since $V^{J^{\star}}_{\alpha}\leq V^{J}_{\alpha}$ for all $\alpha$ and all alternative $(J,\rho)$:
$H(J^{\star},\rho^{\star})=\int_{0}^{1}\!\log V^{J^{\star}}_{\alpha}\,d\alpha\leq\int_{0}^{1}\!\log V^{J}_{\alpha}\,d\alpha=H(J,\rho).$ (16)
@@
Step 4: No alternative possibility assignment does better. Suppose we keep $J^{\star}$ but replace $\rho^{\star}$ with some $\rho^{\prime}$. By Lemma 3.1(ii), any assignment placing higher possibility on a higher-$q$ hypothesis increases the volume of at least one $\alpha$-cut and thus increases $H_{\pi}$. The assignment $\rho^{\star}$ is the unique minimizer among all possibility assignments on $J^{\star}$ because ordering by evidence consistency simultaneously minimizes all $\alpha$-cut volumes.
Step 5 (Race-to-bottom): The prior must not influence selection. Any selection rule incorporating $\pi^{(i)}_{k|k-1}$ into the survivor ranking may retain a high-$q$ hypothesis at the expense of a low-$q$ hypothesis. By Step 2, retaining a higher-$q$ hypothesis increases $V^{J}_{\alpha}$ at the corresponding $\alpha$-level, increasing $H_{\pi}$.
@@
Over multiple steps, prior-based selection compounds: a hypothesis that accumulates high prior possibility may be repeatedly protected from evidence-based elimination, concentrating the possibility distribution on hypotheses that are systematically less evidence-consistent. This is the race-to-bottom failure mode—the possibility gradient is driven by self-reinforcing prior protection rather than evidence. The evidence-only class excludes this failure mode by construction. No rule within the evidence-only class that differs from minimum-$q$ achieves lower $H_{\pi}$.
@@
Uniqueness. For a generic point cloud (no ties in $q^{(i)}_{k}$, holding almost surely for continuous dynamics), $J^{\star}$ is unique and $\rho^{\star}$ is determined pointwise by $\mathrm{Comp}^{(i)}_{k}$.
Nonlinear case. For nonlinear $h$, the innovation-state isotropy condition
$\operatorname{Var}_{x\sim\mathcal{U}(S_{k|k-1})}\!\left\|L_{e}^{-1}(y_{k}-h(x))\right\|^{2}_{2}\leq\operatorname{Var}_{x\sim\mathcal{U}(S_{k|k-1})}\!\left\|L_{\mathrm{core}}^{-1}(x-\bar{x})\right\|^{2}_{2}$ (17)
@@
ensures that the ordering of hypotheses by $q^{(i)}_{k}$ is a consistent proxy for ordering by distance to the evidence-consistent region. Under this condition Steps 2–4 hold for the nonlinear case via first-order Taylor approximation. The condition can fail when the measurement geometry is strongly anisotropic relative to the state-space geometry—for example, when a single range measurement is made from a nearly tangential viewing angle, so that large state displacements along the line of sight produce small innovation changes. Conjecture 7.1 addresses this for the orbital mechanics setting. ∎
@@ 4 Main Theorem
In the falsification regime, subject to the coverage controller operating at PCRB-implied cardinality (i.e., $N_{\mathrm{target}}=\left\lfloor(1-I_{k})\cdot M\right\rfloor$) and $\sigma_{k}$ at steady state:
@@
    The ESPF achieves (18) with equality.
@@
    No other filter in $\mathcal{F}$ achieves lower $H_{\pi,k|k}$ while satisfying (18).
@@
(i) ESPF achieves the PCRB. By Lemma 3.6, the ESPF’s joint strategy $(J^{\star},\rho^{\star})$ minimizes $V^{J}_{\alpha}$ at every $\alpha$-level. The coverage controller sets $N_{\mathrm{target}}=\left\lfloor(1-I_{k})\cdot M\right\rfloor$ survivors, retaining the fraction of the support that the evidence permits per the PCRB. At controller steady state, the integrated log-volume satisfies
$H_{\pi,k|k}=H_{\pi,k|k-1}+\tfrac{n}{2}\log(1-I_{k}),$ (19)
achieving equality in (18). (Existence of steady state follows from the bounded, stationary nature of the Smolyak grid volume and the $\sigma_{k}$ controller; see companion paper [3] for the $\sigma_{k}$ convergence argument.)
(ii) No other filter does better. By the PCRB (Lemma 3.3), every $F^{\prime}\in\mathcal{F}$ satisfies $H_{\pi^{\prime},k|k}\geq H_{\pi,k|k-1}+\frac{n}{2}\log(1-I_{k})$. Since the ESPF achieves equality, $H_{\pi^{\prime},k|k}\geq H^{\mathrm{ESPF}}_{\pi,k|k}$ for all $F^{\prime}\in\mathcal{F}$.
(iii) Uniqueness. By Lemma 3.6, under the isotropy condition the ESPF’s choice $(J^{\star},\rho^{\star})$ is the unique minimizer of $H(J,\rho)$ almost surely. Any alternative filter achieving equality in the PCRB must implement the same selection and possibility assignment; i.e., it must be the ESPF. Any filter using a different rule achieves $H_{\pi^{\prime},k|k}>H^{\mathrm{ESPF}}_{\pi,k|k}$ strictly. ∎
{summarybox}
@@ 4.1 Three Sources of Ignorance, One Optimal Strategy
The $H_{\pi}$ decomposition (8) reveals three distinct aspects of the ESPF’s ignorance management.
@@ 5 The Gaussian Limit: Recovery of the Kalman Filter
In the Gaussian limit, the $\alpha$-cut at level $\alpha$ is the ellipsoid $\{x:(x-\hat{x}_{k})^{\top}\Sigma_{k}^{-1}(x-\hat{x}_{k})\leq-2\log\alpha\}$, with volume $c_{n}(-2\log\alpha)^{n/2}(\det\Sigma_{k})^{1/2}$. Therefore
$H_{\pi}=\frac{1}{2}\log\det\Sigma_{k}+\frac{n}{2}\int_{0}^{1}\!\log(-2\log\alpha)\,d\alpha+\kappa_{n}.$ (20)
The integral $\int_{0}^{1}\log(-2\log\alpha)\,d\alpha=\log 2-\gamma$ is a pure constant ($\gamma$ the Euler–Mascheroni constant). Hence $H_{\pi}=\frac{1}{2}\log\det\Sigma_{k}+\mathrm{const}(n)$, and minimizing $H_{\pi}$ is equivalent to minimizing $\det\Sigma_{k|k}$, which is the Kalman criterion. ∎
Remark 5.3 (Hölder mean structure).
The relationship between the minimax term, $H_{\pi}$, and the Kalman criterion has a unifying interpretation via Hölder means [17]. Applied to the $\alpha$-cut volumes $\{V_{\alpha}\}_{\alpha\in(0,1]}$:
  • $\log\det(\mathrm{MVEE})=\log M_{+\infty}(\{V_{\alpha}\})$: log of the supremum of volumes—the worst-case ignorance (Popperian bound).
  • $H_{\pi}=\log M_{0}(\{V_{\alpha}\})$: log-geometric mean of volumes—integrated ignorance.
  • Kalman’s criterion ($\min\det\Sigma_{k|k}$) == $M_{0}$ in the Gaussian limit.
The Hölder ordering $M_{0}\leq M_{+\infty}$ explains why Claim B ($H_{\pi}$, order 0) is more robustly satisfied than Claim A ($\log\det\mathrm{MVEE}$, order $\infty$) in the numerical validation: the geometric mean is less sensitive to boundary perturbations of a single $\alpha$-cut than the supremum.

### 12 Mar 2026 v2

@@ 3.3 Lemma 3: Minimax Entropy and Evidence-Only Selection
Let $\{\chi^{(i)}\}^{M}_{i=1}$ be the predicted support with whitened squared innovations $q^{(i)}_{k}$, and let $N_{\mathrm{target}}\leq M$ be a fixed survivor count. Consider all pairs $(J,\rho)$ where $J\subseteq\{1,\ldots,M\}$ with $|J|=N_{\mathrm{target}}$ and $\rho:J\to(0,1]$ with $\max_{i\in J}\rho^{(i)}=1$, within the evidence-only comparison class: all $(J,\rho)$ whose selection and ranking depend only on $q^{(i)}_{k}$, not on $\pi^{(i)}_{k|k-1}$.
@@
$H(J,\rho)=\int_{0}^{1}\log V^{J}_{\alpha}\,d\alpha,\quad V^{J}_{\alpha}=c_{n}\bigl(\det\Pi^{J}_{\alpha}\bigr)^{1/2},$ (13)
@@
$J^{*}=\bigl\{i:q^{(i)}_{k}\text{ is among the }N_{\mathrm{target}}\text{ smallest}\bigr\},\quad\rho^{*}(i)=\mathrm{Comp}^{(i)}_{k}=e^{-q^{(i)}_{k}/2},$ (14)
@@
Equivalently, $J^{*}$ is the unique minimizer of $\log\det\bigl(\mathrm{MVEE}(\{\chi^{(i)}:i\in J\})\bigr)$ within the evidence-only class.
@@
We show $(J^{*},\rho^{*})$ minimizes $H(J,\rho)$ by showing it simultaneously minimizes $V^{J}_{\alpha}$ for every $\alpha\in(0,1]$.
Step 1: $\alpha$-cut structure under $\rho^{*}$. Under $\rho^{*}(i)=e^{-q^{(i)}_{k}/2}$, higher possibility corresponds to lower $q^{(i)}_{k}$. For any threshold $\alpha$,
$C_{\alpha}(J^{*},\rho^{*})=\bigl\{i\in J^{*}:e^{-q^{(i)}_{k}/2}\geq\alpha\bigr\}=\bigl\{i\in J^{*}:q^{(i)}_{k}\leq-2\log\alpha\bigr\}.$
At $\alpha\to 0$ this is the full survivor set $J^{*}$.
@@
$q^{(i)}_{k}=\left\|L_{e}^{-1}H(\chi^{(i)}-\chi^{*})\right\|_{2}^{2}+\mathrm{const},$
@@
where $\chi^{*}=H^{\dagger}(y_{k}-b)$ is the evidence-consistent locus. Hence $q^{(i)}_{k}$ measures each hypothesis’s distance to $\chi^{*}$ in the innovation-whitened metric.
@@
Among all $N_{\mathrm{target}}$-element subsets of $\{\chi^{(i)}\}^{M}_{i=1}$, the set geometrically closest to $\chi^{*}$ has the smallest MVEE. This follows from the convexity of $\log\det(\cdot)$ on the positive-definite cone [10]: for any set $J$ containing a hypothesis $j$ with $q^{(j)}_{k}>q^{(i^{*})}_{k}$ for some $i^{*}\notin J$, replacing $j$ with $i^{*}$ (moving the set closer to $\chi^{*}$) does not increase the MVEE determinant, and strictly decreases it for generic (non-degenerate) point clouds.
@@
Therefore $J^{*}$ minimizes $\log\det(\mathrm{MVEE}(\{\chi^{(i)}:i\in J\}))$ among all $N_{\mathrm{target}}$-element subsets. The same reasoning applies at every $\alpha$-level: for each cardinality $k_{\alpha}$, the $k_{\alpha}$ smallest-$q$ hypotheses minimize $V^{J}_{\alpha}$. Hence $C_{\alpha}(J^{*},\rho^{*})$ minimizes $V^{J}_{\alpha}$ for every $\alpha\in(0,1]$.
@@
Step 3: Pointwise minimum volume implies minimum $H_{\pi}$. Since $V^{J^{*}}_{\alpha}\leq V^{J}_{\alpha}$ for all $\alpha$ and all alternative $(J,\rho)$:
$H(J^{*},\rho^{*})=\int_{0}^{1}\log V^{J^{*}}_{\alpha}\,d\alpha\leq\int_{0}^{1}\log V^{J}_{\alpha}\,d\alpha=H(J,\rho).$
@@
Step 4: No alternative possibility assignment does better. Suppose we keep $J^{*}$ but replace $\rho^{*}$ with some $\rho^{\prime}$. By the monotone concentration property of $H_{\pi}$, any assignment placing higher possibility on a higher-$q$ hypothesis increases the volume of at least one $\alpha$-cut and thus increases $H_{\pi}$. The assignment $\rho^{*}$ is the unique minimizer among all possibility assignments on $J^{*}$ because ordering by evidence consistency simultaneously minimizes all $\alpha$-cut volumes.
@@
Step 5 (Race-to-bottom): The prior must not influence selection. Any selection rule incorporating $\pi^{(i)}_{k|k-1}$ into the survivor ranking may retain a high-$q$ hypothesis at the expense of a low-$q$ hypothesis. By Step 2, retaining a higher-$q$ hypothesis increases $V^{J}_{\alpha}$ at the corresponding $\alpha$-level, increasing $H_{\pi}$. Over multiple steps, prior-based selection compounds: a hypothesis that accumulates high prior possibility may be repeatedly protected from evidence-based elimination, concentrating the possibility distribution on hypotheses that are systematically less evidence-consistent. This is the race-to-bottom failure mode—the possibility gradient is driven by self-reinforcing prior protection rather than evidence. The evidence-only class excludes this failure mode by construction. No rule within the evidence-only class that differs from minimum-$q$ achieves lower $H_{\pi}$.
@@
Uniqueness. For a generic point cloud (no ties in $q^{(i)}_{k}$, holding almost surely for continuous dynamics), $J^{*}$ is unique and $\rho^{*}$ is determined pointwise by $\mathrm{Comp}^{(i)}_{k}$.
@@
Nonlinear case. For nonlinear $h$, condition (10) ensures that the ordering of hypotheses by $q^{(i)}_{k}$ is a consistent proxy for ordering by distance to the evidence-consistent region. Under this condition Steps 2–4 hold for the nonlinear case via first-order Taylor approximation. The condition can fail when the measurement geometry is strongly anisotropic relative to the state-space geometry—for example, when a single range measurement is made from a nearly tangential viewing angle, so that large state displacements along the line of sight produce small innovation changes. Proposition 3.10 establishes this condition for the orbital mechanics setting. ∎
@@
Proposition 3.10 (Isotropy for orbital mechanics).
Let the state be the 7-dimensional Cartesian position-velocity-drag vector $x=(r,\dot{r},C_{d})\in\mathbb{R}^{7}$, propagated under the two-body problem or its $J_{2}$-perturbed extension. Let the measurement model be range $\rho(x)=\|r-r_{s}\|$ and range-rate $\dot{\rho}(x)=(r-r_{s})\cdot\dot{r}/\rho(x)$, where $r_{s}$ is the ground station position. Let the observation interval satisfy $\Delta t\ll T/4$, where $T$ is the orbital period. Then the innovation-state isotropy condition (10) holds to first order in the linearization residual, except in a measure-zero set of viewing geometries in which the line-of-sight vector $(r-r_{s})/\rho$ is orthogonal to the velocity $\dot{r}$ (tangential viewing).
Proof sketch.
Let $H=\partial h/\partial x|_{\bar{x}}$ be the measurement Jacobian evaluated at the support centroid. The whitened innovation variance decomposes as:
$\mathrm{Var}\!\left[\left\|L_{e}^{-1}(y-h(x))\right\|^{2}\right]\approx\mathrm{Var}\!\left[\left\|L_{e}^{-1}H(x-\bar{x})\right\|^{2}\right]+O(\|\delta H\|^{2}),$
where $\delta H$ is the linearization residual. The first term equals $\mathrm{tr}(L_{e}^{-1}H\Pi_{\mathrm{core}}H^{\top}L_{e}^{-\top})$, which by the Cauchy–Schwarz inequality on the Frobenius inner product satisfies:
$\mathrm{tr}(L_{e}^{-1}H\Pi_{\mathrm{core}}H^{\top}L_{e}^{-\top})\leq\|L_{e}^{-1}H\|_{F}^{2}\cdot\lambda_{\max}(\Pi_{\mathrm{core}}).$
For range and range-rate measurements in LEO, $H$ has rank 2 and $\|L_{e}^{-1}H\|_{F}^{2}$ is bounded by $\sigma_{R}^{-2}+\sigma_{\dot{R}}^{-2}$, which under nominal measurement noise ($\sigma_{R}=1$ m, $\sigma_{\dot{R}}=10^{-5}$ km/s) and a support cloud spanning tens of kilometers satisfies condition (10) with margin. The bound degrades as $(r-r_{s})/\rho\to\dot{r}/\|\dot{r}\|$ (tangential geometry), at which point $\partial\dot{\rho}/\partial r$ becomes nearly parallel to $\partial\rho/\partial r$ and the rank-2 Jacobian loses its geometric leverage—the failure mode identified in Lemma 3.7. When this occurs, the EWM surprisal indicator escalates in real time, providing an operational signal that the isotropy condition is under stress and that the filter’s Popperian updates should be interpreted with reduced confidence; the filter thus degrades gracefully rather than silently. For $\Delta t\ll T/4$, the linearization residual $O(\|\delta H\|^{2})$ remains small relative to the leading term, completing the sketch. ∎
@@ 4 Main Theorem
In the falsification regime, subject to the coverage controller operating at PCRB-implied cardinality (i.e., $N_{\mathrm{target}}=\lfloor(1-I_{k})\cdot M\rfloor$) and $\sigma_{k}$ at steady state:
@@
    The ESPF achieves the PCRB with equality.
@@
    No other filter in $\mathcal{F}$ achieves lower $H_{\pi,k|k}$ while satisfying the PCRB.
@@
(i) ESPF achieves the PCRB. By Lemma 3.7, the ESPF’s joint strategy $(J^{*},\rho^{*})$ minimizes $V^{J}_{\alpha}$ at every $\alpha$-level. The coverage controller sets $N_{\mathrm{target}}=\lfloor(1-I_{k})\cdot M\rfloor$ survivors, retaining the fraction of the support that the evidence permits per the PCRB. At controller steady state, the integrated log-volume satisfies
$H_{\pi,k|k}=H_{\pi,k|k-1}+\frac{n}{2}\log(1-I_{k}),$
achieving equality in the PCRB.
(ii) No other filter does better. By the PCRB (Lemma 2), every $F^{\prime}\in\mathcal{F}$ satisfies $H^{\prime}_{\pi,k|k}\geq H_{\pi,k|k-1}+\frac{n}{2}\log(1-I_{k})$. Since the ESPF achieves equality, $H^{\prime}_{\pi,k|k}\geq H^{\mathrm{ESPF}}_{\pi,k|k}$ for all $F^{\prime}\in\mathcal{F}$.
(iii) Uniqueness. By Lemma 3.7, under the isotropy condition the ESPF’s choice $(J^{*},\rho^{*})$ is the unique minimizer of $H(J,\rho)$ almost surely. Any alternative filter achieving equality in the PCRB must implement the same selection and possibility assignment; i.e., it must be the ESPF. Any filter using a different rule achieves $H^{\prime}_{\pi,k|k}>H^{\mathrm{ESPF}}_{\pi,k|k}$ strictly. ∎
@@ 4.1 Three Sources of Ignorance, One Optimal Strategy
The $H_{\pi}$ decomposition reveals three distinct aspects of the ESPF’s ignorance management.
@@ 5 The Gaussian Limit: Recovery of the Kalman Filter
In the Gaussian limit, the $\alpha$-cut at level $\alpha$ is the confidence ellipsoid $\mathcal{E}_{\alpha}=\{x:(x-\hat{x}_{k})^{\top}\Sigma_{k}^{-1}(x-\hat{x}_{k})\leq-2\log\alpha\}$, with volume $V_{\alpha}=c_{n}(-2\log\alpha)^{n/2}(\det\Sigma_{k})^{1/2}$. Substituting into $H_{\pi}=\int_{0}^{1}\log V_{\alpha}\,d\alpha$ and evaluating the constant integral $\int_{0}^{1}\log(-2\log\alpha)\,d\alpha=\log 2-\gamma$ gives $H_{\pi}=\frac{1}{2}\log\det\Sigma_{k}+\mathrm{const}(n)$. Minimizing $H_{\pi}$ is therefore equivalent to minimizing $\det\Sigma_{k|k}$, which is the Kalman criterion. Remark 5.3 derives this identity from the Hölder mean structure and shows it is the Gaussian specialization of the general possibilistic functional. ∎
Remark 5.3 (Hölder mean structure and the probability–possibility bridge).
The continuous Hölder power mean of the family $\{V_{\alpha}\}_{\alpha\in(0,1]}$ at order $p$ is
$\log M_{p}\bigl(\{V_{\alpha}\}\bigr)=\frac{1}{p}\log\int_{0}^{1}V_{\alpha}^{p}\,d\alpha,\quad p\neq 0,$ (16)
with the geometric mean limit
$\log M_{0}\bigl(\{V_{\alpha}\}\bigr)=\lim_{p\to 0}\,\frac{1}{p}\log\int_{0}^{1}V_{\alpha}^{p}\,d\alpha=\int_{0}^{1}\log V_{\alpha}\,d\alpha,$ (17)
and the extremes $\log M_{+\infty}=\log\sup_{\alpha}V_{\alpha}$ and $\log M_{-\infty}=\log\inf_{\alpha}V_{\alpha}$ as $p\to\pm\infty$. The three filter criteria emerge as evaluations of this single functional at different orders:
  • $p\to+\infty$: $\log M_{+\infty}=\log V_{0}=\log\det(\mathrm{MVEE})$ — the outermost $\alpha$-cut volume, worst-case ignorance, the Popperian minimax bound operative in the falsification regime.
  • $p\to 0$: $\log M_{0}=\int_{0}^{1}\log V_{\alpha}\,d\alpha=H_{\pi}$ — the log-geometric mean of $\alpha$-cut volumes, integrated possibilistic ignorance, the ESPF optimality criterion.
  • $p\to$-$\infty$: $\log M_{-\infty}=\log V_{1}$ — the innermost $\alpha$-cut volume (the anchor hypothesis neighborhood), best-case ignorance, a lower floor.
The Hölder ordering $M_{-\infty}\leq M_{0}\leq M_{+\infty}$ (strict for non-constant $\{V_{\alpha}\}$) is the mathematical expression of the epistemological hierarchy: the ESPF’s integrated ignorance is bounded above by the Popperian worst-case and below by the best-case floor.
The probability–possibility bridge. The same functional $\log M_{0}(\{V_{\alpha}\})$ recovers the Kalman criterion when evaluated under Gaussian geometry. In the possibilistic case, $V_{\alpha}$ is the MVEE volume of hypotheses with $\pi^{(i)}\geq\alpha$. In the Gaussian case, the natural analog of the $\alpha$-cut is the confidence ellipsoid
$\mathcal{E}_{\alpha}=\bigl\{x:(x-\hat{x})^{\top}\Sigma^{-1}(x-\hat{x})\leq-2\log\alpha\bigr\},$
with volume $V_{\alpha}^{\mathrm{Gauss}}=c_{n}(-2\log\alpha)^{n/2}(\det\Sigma)^{1/2}$. Substituting into $\log M_{0}$:
$H_{\pi}^{\mathrm{Gauss}}$ $=\int_{0}^{1}\log V_{\alpha}^{\mathrm{Gauss}}\,d\alpha$
$=\frac{1}{2}\log\det\Sigma+\frac{n}{2}\int_{0}^{1}\log(-2\log\alpha)\,d\alpha+\kappa_{n}$
$=\frac{1}{2}\log\det\Sigma+\mathrm{const}(n),$ (18)
recovering the Kalman criterion exactly, since $\int_{0}^{1}\log(-2\log\alpha)\,d\alpha=\log 2-\gamma$ is a pure constant. Possibility theory and probability theory are not competing frameworks for this class of problems: they are evaluations of the same ignorance functional $\log M_{0}(\{V_{\alpha}\})$ under different geometries of $\alpha$-cuts. The possibilistic $\alpha$-cut is defined by the MVEE of a possibility level set; the probabilistic $\alpha$-cut is defined by a Gaussian confidence ellipsoid. The functional that integrates their log-volumes is identical. The Kalman filter is the Gaussian specialization of the ESPF’s optimality criterion, not a separate invention.